OpenStudy (anonymous):

A ship leaves a pier on a bearing of N50degreeW and travels for 75miles. It turns and sails on a bearing of S 20degree W for 50miles. How far from the pier is the ship?

OpenStudy (anonymous):

Question isnt spelled right its Notical Miles lol :P

OpenStudy (anonymous):

But! I guess its like way over 75+50 miles away from that

OpenStudy (anonymous):

oh but do you know how to solve this?

OpenStudy (anonymous):

Let me ask my team Srry went to eat Mcdonalds Woot!

OpenStudy (dumbcow):

 we can solve for x by using law of cosines: The pier, boat, and point where it turned make a triangle of lengths, 75,50,x angle opposite of side x is 70 degrees --> x^2 = 75^2+50^2 -2(75)(50)cos(70) x = 74.564 miles